OpenStudy (anonymous):

y=ln((x^2 sinx)/sqrt(1+x)) calculate the derivative in 3 different ways any help for this exercise?

OpenStudy (kc_kennylau):

\[\Large y=\ln\left(\frac{x^2\sin x}{\sqrt{1+x}}\right)\]

OpenStudy (kc_kennylau):

Don't tell me you're going to use the first principle

OpenStudy (kc_kennylau):

I don't think I can give you 3 different methods, but I think I can give you one method

OpenStudy (kc_kennylau):

Chain rule+quotient rule+product rule+chain rule

OpenStudy (anonymous):

this is the only way to sovle this exercise?

OpenStudy (kc_kennylau):

I'm not sure because one Mathematics question can have many approaches but only one solution

OpenStudy (kc_kennylau):

This is perhaps the beauty of Mathematics

OpenStudy (kc_kennylau):

The first principle may be feasible

OpenStudy (anonymous):

1) u substitution 2) begin with a chain rule and go from there

OpenStudy (kc_kennylau):

But the arithmetic would be tedious

OpenStudy (anonymous):

could one of you perhaps help me with my question too? http://openstudy.com/study#/updates/5341dc55e4b09298524f057f